A sample of 1600 computer chips revealed that 26% of the chips fail in the first 1000 hours of their use. The company's promotional literature states that 24% of the chips fail in the first 1000 hours of their use. The quality control manager wants to test the claim that the actual percentage that fail is different from the stated percentage. Find the value of the test statistic. Round your answer to two decimal places.
Answer #1

H0: Null Hypothesis: P = 0.24

HA: Alternative Hypothesis: P eq 0.24

n = Sample Size = 1600

p = Sample proportion = 0.26

P = Population Proportion = 0.24

Q = 1 - P =0.76

SE = sqrt{PQ/n}

= V/0.24 × 0.76/1600= 0.0107

Test staistic is:

Z= (0.26 - 0.24)/0.0107 = 1.87

So,

Answer is:

1.87